# Discretized Quantum Exhaustive Search for Variational Quantum Algorithms

### v1

@@ 3.3 Transverse-Field Ising Problem
In Figures  9 and  10, we show results of energy landscaping of such Hamiltonians with 3 qubits. The landscaping is done using a full MUB set. We colored the results from different bases and show that by having different parameters for the Hamiltonian, the basis that has a state with the lowest energy is different. Because of the locality of each term in the sum of operators that create the Hamiltonian, MUB states from high dimension can be traced out to dimension 2 and their expectation value can be calculated exactly using a classical computer. Such calculation can be done only if the state is known and can be represented directly as a quantum state and not only as a unitary transformation that evolves some initial state (which in that case such transformation should be calculated first, and in case of a very high order, might not be possible to calculate on a classical computer). For that reason, such a calculation can’t be used during the optimization process after using a MUB state as the initial state for an ansatz.
@@
Figure 9: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the computational basis (blue color). The Hamiltonian coefficients are: $H=0.04645122\cdot ZZI+0.04645122\cdot IZZ+0.27498273\cdot XII+0.27498273\cdot IXI+0.27498273\cdot IIX$
@@
Figure 10: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the Hadamard basis (orange color). The Hamiltonian coefficients are: $H=0.61436456\cdot ZZI+0.61436456\cdot IZZ+0.32435029\cdot XII+0.32435029\cdot IXI+0.32435029\cdot IIX$
3.4 Combinatorial Problems
@@
We used a full set of MUB states of 3 qubits for the efficient partial DQES method to explore the landscape of the Max-Cut problem (additional details in the appendix). We used 3 qubits MUB states, and kept the rest of the qubits at the zero state. For each experiment, we tested all possible triples of qubit locations to be defined as the qubits of the MUB state.
@@
As an example, we used a random graph with 8 nodes, and sampled it’s energy landscape using partial-MUB states. We can see from the spread of the different results, that the cost function has the most differences in values in the computational basis, while in the other bases, most of the states yield the same result. Such knowledge can be used to limit the ansatz to specific states in the Hilbert space of the problem. We also tried to find the optimal solution by running an optimization process using the VQE algorithm (using COBYLA as optimization algorithm). We compared the results between different initial state - 2 MUB states which had low energy in the energy landscaping step, and 2 random states which had low energy relative to other random states that were randomized (we sampled the same amount of random states and MUB states). The results are shown in Fig.11
@@
Figure 11: Max-Cut problem example. (a) Exhaustive search of all the possible tensor products of 3 qubit MUB states with the rest of the qubits in the zero state, and their cost function value for a random 8 nodes Max-Cut problem. (b) Convergence graphs of VQE runs on the 8 nodes Max-Cut problem, starting with different initial guesses - 2 choices of MUB states and 2 random states.
4 Conclusion
@@
In this paper, we have tried to tackle the most dominant problems, local minima and barren plateaus, of one of the promising algorithmic approaches to the NISQ era of quantum computing - VQAs.
@@
We have developed a method to efficiently sample the Hilbert space of the cost function of a VQA problem in a efficient partial DQES approach. We used MUB states in order to extract maximum information out from such sampling, in an ansatz agnostic method. We tested our method on multiple different problems, including different molecular electronic structure problems, Transverse-Field Ising model problems, and Max-Cut problems.
@@
We showed that such a method can be used in order to learn characteristics of the landscape of the cost function and some useful information about the given problem at hand, like good approximations of the problem and good initial guesses to the solution. We managed to learn good approximations and initial guesses regarding the tested problems, which are similar to those achieved by experts in the problems’ fields. We have extracted such information without using prior expert knowledge. We believe that the information learned on the cost function landscape can be used to help avoid barren plateaus during the optimization process done in a VQA.
4.1 Open problems for Future research
@@
Our research introduces new ansatz and methods to deal with the trainability of VQAs, but further research is still needed in order to make these algorithms achieve results better than the best classical computers using today’s available quantum hardware. Such further research might include -
@@
  • Try to use the information gained from the DQES search or the efficient partial DQES to avoid BP and improve the algorithm convergence rate or final accuracy.
  • Apply DQES and partial DQES to problems with larger number of qubits.

### v2

@@ III.3 Transverse-Field Ising Problem
In Figures 8 and 9, we show results of energy landscape calculation of such Hamiltonians with 3 qubits. The landscape calculation is done using a full MUB set. We colored the results from different bases and showed that by having different parameters for the Hamiltonian, the basis that has a state with the lowest energy is different. Because of the locality of each term in the sum of operators that create the Hamiltonian, MUB states from high dimension can be traced out to dimension 2 and their expectation value can be calculated exactly using a classical computer. Such calculation can be done only if the state is known and can be represented directly as a state-vector and not only as a unitary transformation that evolves some initial state (which in that case such transformation should be calculated first, and in case of a very high order, might not be possible to calculate on a classical computer). For that reason, such a calculation can’t be used during the optimization process after using a MUB state as the initial state for an ansatz.
@@
Figure 8: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the computational basis (blue color). The Hamiltonian coefficients are: $\mathbb{H}=0.04645122\cdot ZZI+0.04645122\cdot IZZ+0.27498273\cdot XII+0.27498273\cdot IXI+0.27498273\cdot IIX$
@@
Figure 9: Energy landscape of a transverse-field Ising Hamiltonian. each point represents a different state, while all of the states from the same basis have the same color. The lowest energy state is part of the Hadamard basis (orange color). The Hamiltonian coefficients are: $\mathbb{H}=0.61436456\cdot ZZI+0.61436456\cdot IZZ+0.32435029\cdot XII+0.32435029\cdot IXI+0.32435029\cdot IIX$
III.4 Combinatorial Problems
@@
We used the efficient partial DQES method with a full set of MUB states of 3 qubits to explore the landscape of the Max-Cut problem [7] (more details in appendix E). We used 3-qubit MUB states, and kept the rest of the qubits at the $|0\rangle$ state. For each experiment, we tested all possible triples of qubit locations to be defined as the qubits of the MUB state.
@@
For our example, we used a random graph with 8 nodes, and sampled its energy landscape using partial-MUB states. We can see from the spread of the different results, that the cost function has the most differences in values in the computational basis. In the other bases, in contrast, most of the states yield the same result. Such knowledge can be used to limit the ansatz to specific states in the Hilbert space of the problem. We also tried to find the optimal solution by running an optimization process using the VQE algorithm (using COBYLA as optimization algorithm and hardware efficient ansat with 3 layers and linear entanglement). We compared the results between different initial states: 2 MUB states which had low energy in the energy landscape calculation step, and 2 random states which had low energy relative to other random states that were randomized (we sampled the same amount of random states and MUB states). To create the random states we sampled vectors of parameters for the ansatz. For each parameterized gate we randomized a rotation angle uniformly from the range $0,2\pi)$. For each such random parameters vector we calculated the cost function for the problem’s graph. We than ran the VQE optimization process twice, using the parameters from the two vectors which resulted in the lowest cost as initial parameters for the optimization. The results are shown in Fig. 10
@@
Figure 10: Max-Cut problem example. (a) Exhaustive search of all the possible tensor products of 3 qubit MUB states with the rest of the qubits in the $|0\rangle$ state, and their cost function value for a random 8 nodes Max-Cut problem. The points of all the states from each basis of the MUB from all possible qubit allocations share the same background color. The ordering of the points for each MUB is the same as described as the beginning of section III. 8 qubits were used to encode this problem and 56 qubit allocations were tested. For each qubit allocation, the energy for 72 different MUB states was calculated, a total of 4032 points calculated. (b) Convergence graphs of VQE runs on the 8 nodes Max-Cut problem, starting with different initial guesses - 2 choices of MUB states and 2 random states.
IV Conclusion
@@
In this paper, we tried to tackle local minima and barren plateaus, which are the most dominant problems of VQAs — one of the promising algorithmic approaches to the NISQ era of quantum computing.
@@
We have developed a method to efficiently sample the Hilbert space of the cost function of a VQA problem, using the “efficient partial DQES” approach. We used MUB states in order to extract maxinmal information out from such sampling, in an ansatz-agnostic method. We tested our method on multiple different problems, including different molecular electronic structure problems, Transverse-Field Ising model problems, and Max-Cut problems.
@@
We showed that such a method can be used in order to learn characteristics of the landscape of the cost function and some useful information about the given problem at hand, such as good approximations of the problem and good initial guesses to the solution. We managed to learn good approximations and initial guesses regarding the tested problems, which are similar to those achieved by experts in the problems’ fields. We have extracted such information without using prior expert knowledge. We believe that the information learned on the cost function landscape can be used to help avoid barren plateaus during the optimization process done in a VQA.
IV.1 Open problems for Future research
@@
Our research introduces new ansatzes and methods to deal with the convergence issues of VQAs, but further research is still needed in order to make these algorithms achieve results better than the best classical computers using today’s available quantum hardware. Such further research might include -
@@
  • In contrast with the trivial choice of the $|{0}\rangle$ state for non-MUB qubits in efficient partial DQES, use a random choice (from a pre-determined set) for the single-qubit states and explore the implications for different problem types.
  • Try to use the information gained from the DQES search or the efficient partial DQES to avoid barren plateaus and local minima and improve the algorithm convergence rate or final accuracy.
  • Apply DQES and partial DQES to problems with larger numbers of qubits.
@@
  • Further investigate whether the efficient partial DQES does better than random states for NP problems as seen in two examples we provided in Fig. 10.
  • Further investigate advantages and disadvantages of DQES and efficient partial DQES when compared to random initial states of the same amount, for problems beyond NP, and for problems in BQP (such as variational quantum factoring).
@@ Acknowledgements:
Appendix A Variational quantum algorithms
Variational quantum algorithms represent a pioneering approach within the field of quantum computing, developed to tackle the challenges that arise in the NISQ era and offer a practical bridge between current limited quantum hardware and problem-solving [14]. At their core, these algorithms employ quantum circuits with tunable parameters, often referred to as variational circuits or parametric circuits, to address a wide range of computational tasks. By iteratively adjusting these parameters and leveraging classical optimization techniques, variational quantum algorithms seek to find the optimal configuration that minimizes a specific objective function. This hybrid approach combines the power of quantum superposition and entanglement with classical optimization’s robustness, making it suitable for tackling problems in fields as diverse as quantum chemistry, optimization, machine learning, and materials science.
The various variational quantum algorithms are defined mainly by their cost function, and the structure of the associated quantum circuit which allows the efficient evaluation of that cost function. The efficiency is measured in the ability to calculate the cost function using only a polynomial (in the problem size) number of circuit executions at each iteration of the classical optimization process.
The evaluation of the cost function is done by calculating the expectation values of some operators derived from the cost function. These expectation values are calculated in a state $|\psi(\vec{\theta})\rangle$, which is a parameterized state prepared by the used ansatz.
$C(\vec{\theta})=\sum_{k}f_{k}(\langle\psi(\vec{\theta})|O_{k}|\psi(\vec{\theta})\rangle)$ (6)
Where $f_{k}$ is a set of functions, $O_{k}$ is the set of the operators that are calculated on the state prepared by the ansatz.
At each iteration, a classical computer calculates the next parameter vector that will be used in the ansatz parameters. The quantum computer repeatedly measures the operators $O_{k}$ over the parametric state. A classical computer evaluates the functions $f_{k}$ and constructs the value of the cost function $C$. The new evaluation is fed into the optimization algorithm, which outputs the next set of parameters. This process continues until a convergence or other stop criteria (like a maximum number of iterations) is reached. The goal of this process is to find an optimal parameters vector:
$\vec{\theta}_{opt}=\arg\min_{\vec{\theta}}C(\vec{\theta})$ (7)
This iterative process has several benefits that are important in the NISQ era. Among those benefits is the ability to break down a deep circuit into a batch of shallower circuits, which in turn allows to calculate the result with less noise. Another important benefit is the partial resilience to coherent errors that arise from this iterative process [36, 42]. This resilience is related mainly to the ansatz part, and is a result of the fact that the prepared state is changing at each iteration to a state that gaints a better value from the cost function, fixing coherent errors by using new parameters. For example, if the computer has an over-rotation error for one of its gates as a result of a miscalibrated gate, the parameters vector can shift slightly in the following iterations to account for this over-rotation. VQAs still have many challenges, as achieving the global optimum of ansatz parameters can be computationally intensive, and the landscape of the objective function can feature many local minima. In many cases, as the problem size increases, the gradients of the cost function become exponentially small, making the search for the optimal solution almost impossible. This challenge is referred to as the barren plateaus problem [35].
Appendix B Ansatzes
An ansatz is typically constructed as a quantum circuit with a specific structure, such as a sequence of quantum gates, where some or all of the gates have adjustable parameters. These parameters are the variables that can be optimized to find the best approximation to the target quantum state, which in turn optimizes the algorithm’s cost function. The ansatz prepares a parameterized quantum state -
$|\psi(\vec{\theta})\rangle=U(\vec{\theta})|\psi_{0}\rangle$ (8)
while $|\psi(\vec{\theta})\rangle$ is the result parameterized quantum state, $U(\vec{\theta})$ is the parameterized quantum circuit (the ansatz) and $|\psi_{0}\rangle$ is some initial quantum state.
The term ”ansatz” is derived from the German word for ”assumption” or ”postulate,” and it reflects the idea that you are making an educated guess about the form of the quantum state you want to prepare.
Many VQAs can evaluate their cost function using various ansatzes, each such ansatz having its advantages and disadvantages. The choice of an ansatz is a critical design decision in VQAs. It influences the algorithm’s convergence speed, accuracy, and suitability for a given problem. It also determines the quantum circuit’s expressiveness - its ability to represent complex quantum states or solutions to optimization problems. Different ansatz structures are suited to different types of problems and may have varying levels of flexibility and computational efficiency.
For many ansatz designs, there is a trade-off between the quantum circuit’s expressiveness, the depth of the quantum circuit, and the number of parameters associated with it, which might influence its convergence rate.
In some cases, domain-specific ansatz circuits are designed with prior knowledge about the problem at hand. These ansatzes are usually referred to as problem-inspired ansatzes. For instance, in quantum chemistry applications, one might use an ansatz tailored to represent the electronic structure of a molecule. An example of such a tailored ansatz for the electronic structure of a molecule is the Unitary Coupled Cluster (UCC) ansatz. Problem-inspired ansatzes usually have relatively small search space (only between states that are meaningful to the problem) which allows for fast convergence of the algorithm. On the other hand, in many cases this comes with the disadvantage of a deep quantum circuit, creating a strong noise that makes the results inaccurate.
Another common type of ansatzes is the hardware-efficient ansatz. These ansatzes are generic and can be used in multiple problems and algorithms. The structure of these ansatzes is designed to minimize the resource requirements and circuit depth of the ansatz, while still providing a reasonably expressive representation of quantum states. This structure usually also takes into account the native gates set of the device and avoids the use of complex multi-qubit gates. By that, complex gate decompositions are avoided, which can introduce additional errors and increase the gate depth. The term ”hardware-efficient” underscores the aim of creating quantum circuits that can be implemented efficiently on existing or near-term quantum hardware, which typically has limitations in terms of qubit connectivity, gate fidelities, and gate times.
Appendix C Variational Quantum Eigensolver (VQE)
The Variational Quantum Eigensolver (VQE) [44, 53] is a variational algorithm whose goal is to find the smallest eigenvalue of a given Hermitian matrix (usually described as a Hamiltonian of a system).
$\mathbb{H}|\psi\rangle=E_{0}|\psi\rangle$ (9)
The algorithm finds $E_{0}$, which is usually referred to as the ground state energy of the given Hamiltonian.
The algorithm’s input is a decomposition of the Hermitian matrix into a polynomial number of measurable observables (usually tensored Pauli operators), a chosen ansatz, and an initial guess for the initialization of the parameters. The input Hamiltonian is represented as -
$\mathbb{H}=\sum_{i}^{L}h_{i}O_{i}$ (10)
where $L$ is a polynomial number in the number of qubits $L=poly(n)$, $O_{i}$ are measurable observables, each with a coefficient $h_{i}$ (each observable can act on all or some of the qubits).
The algorithm uses the variational principle, which states that the expectation value of a Hamiltonian, calculated in any state $|\psi\rangle$, is greater than or equal to the ground state energy of that Hamiltonian:
$\langle\psi|\mathbb{H}|\psi\rangle\geq E_{0}$ (11)
The cost function of the algorithm is straightforward: the expectation value of the problem’s Hamiltonian in the parameterized state created by the ansatz
$C(\vec{\theta})=\langle\psi(\vec{\theta})|\mathbb{H}|\psi(\vec{\theta})\rangle$ (12)
This is calculated by the sum of the expectation values in the state created by the ansatz for each measurable observable in the Hamiltonian decomposition:
$C(\vec{\theta})=\sum_{i}^{L}h_{i}\langle\psi(\vec{\theta})|O_{i}|\psi(\vec{\theta})\rangle$ (13)
In the common case of the measurable observables being Pauli operators, measuring their expectation values can be done by adding single qubit gates, before the measurement of the qubits. The Pauli operator can be broken down to a tensor product of single qubit Pauli operators, so the gates to be added for each qubit are simply defined by the local Pauli operator on that qubit. Assuming that the measurement is done in the computational basis, the gates that should be added are as follows - a Hadamard gate for measuring the $X$ operator, a $S^{\dagger}$ gate followed by a Hadamard gate for measuring $Y$ operator. For measuring the $Z$ operator, no gate should be added.
Appendix D Molecular Electronic Structure Problem
One of the most common problems solved using the VQE algorithm is the molecular electronic structure problem. In this problem, an approximation of the ground state energy of the steady state of a molecule should be calculated. The golden standard in computational chemistry for such an approximation is called chemical accuracy and is considered to be the accuracy required to make realistic chemical predictions for some important chemical reactions (for example, the bond distance of a molecule or the energy required to break a molecular bond). This approximation is a fixed (does not scale with problem size) additive approximation equal to 1 [kcal/mol] which is about 1.6 [milli-Hartree/molecule]. Such approximation is made relative to the perfect solution of the problem if one can be calculated analytically, or to the results of a chemical experiment done in a lab in case an analytical solution can not be achieved.
In order to map such problems into a finite space that can be solved on a computer (either classical or quantum), some assumptions and approximations must be used. In our research we used common assumptions and approximations for the electronic structure Hamiltonian:
  • relativistic effects are completely neglected.
  • The Born–Oppenheimer approximation is used.
  • The molecular orbitals are assumed to be a linear combination of a finite number of basis functions (for example, using the STO-nG family of basis functions, while the $n$ is the number of basis functions).
  • Only a finite set of molecular spin orbitals are taken into account in the calculation, while the others are defined before the calculation as either empty or fully occupied (the number of spin orbitals will define the problem size and therefore the runtime complexity of the problem). Each molecular orbital is treated as 2 spin orbitals, therefore the number of relevant $M$ spin-orbitals is always an even number.
The electronic structure Hamiltonian can be written as follows [25, 53]:
$\mathbb{H}=\sum_{i,j}^{M}h_{i,j}a_{i}^{\dagger}a_{j}+\sum_{i,j,k,m}^{M}h_{i,j,k,m}a_{i}^{\dagger}a_{j}^{\dagger}a_{k}a_{m}$ (14)
while $h_{i,j}$ and $h_{i,j,k,m}$ are the coefficients of the single and two electron energy terms, respectively. When summing over spin orbitals instead of electrons we lose in the equation the conservation of the number of electrons, and might enforce this conversation rule to reduce the number of qubits needed to encode the Hamiltonian. The size of this Hamiltonian (if written as a matrix) is exponential in the number of spin orbitals taken into account in the calculation. The $h_{i,j}$ and $h_{i,j,k,m}$ terms can be calculated using the one- and two-body integrals:
$\begin{split}&h_{i,j}=\frac{1}{2}\langle i|\nabla_{i}^{2}|j\rangle+\langle i|\frac{Z_{A}}{r_{iA}}|j\rangle=\\ &\int\varphi_{i}(r)^{*}(-\frac{\hbar^{2}}{2m_{e}}\nabla^{2}-\sum_{A}^{P}\frac{e^{2}}{4\pi\epsilon_{0}}\frac{Z_{A}}{|r_{i}-R_{A}|})\varphi_{j}(r)\,dx\end{split}$ (15)
$\begin{split}&h_{i,j,k,m}=\langle i,j|\frac{1}{r_{ij}}|k,m\rangle=\\ &\frac{e^{2}}{4\pi\epsilon_{0}}\int\frac{\varphi_{i}(r)^{*}\varphi_{j}(r)^{*}\varphi_{k}(r)\varphi_{m}(r)}{|r_{1}-r_{2}|}\,dx_{1}\,dx_{2}\end{split}$ (16)
while $e$ is the elementary charge constant, $m_{e}$ is the electron mass, $Z_{A}$ is the atomic number of the nucleus $A$, $R_{A}$ is the nuclear $A$ position, $r_{i}$ is the electron $i$ position, $\varphi_{i}$ is the electronic wavefunction of the spin orbital $i$ (represented by a linear set of basis functions) and $x_{i}$ is subsuming both the electron position and spin ($x_{i}=(r_{i},\sigma)$).
The coefficients of the Hamiltonian are calculated based on the specific molecule of the problem, the basis functions used, the specific orbitals that are taken into account, and the distance (or coordinates in problems with >1>1 dimensions) between the atoms’ nucleus. Finding the ground state energy of this Hamiltonian is usually referred as the “full configuration interaction” (FCI) problem. The energy of the ground state might be different for each chosen basis functions, orbitals and atoms’ coordinates and will be equal or smaller than the Hartree-Fock energy (as additional approximations are taken in the Hartree-Fock energy case). Solving a similar problem up to a multiplicative error has proven to be a QMA problem [10]. There exists a naive method to solve this problem, by diagonalizing the problem’s Hamiltonian (as described in equation 14) and finding its smallest eigenvalue. As the size of the Hamiltonian is exponential in the number of spin orbitals taken into account in the calculation, this method is inefficient. One can use this method to compare experimental results to the exact FCI result in small-size problems.
Appendix E Combinatorial Problems
The Max-Cut problem is a well-known combinatorial optimization problem with broad applications in fields such as network design, logistics, and computer science. Given an undirected graph, the goal is to partition the graph’s vertices into two disjoint sets (usually denoted as ”cutting” the graph) to maximize the total number of edges crossing the partition. This problem is NP-hard, meaning that finding the optimal partition for large graphs becomes computationally challenging.
The cost function we used for this problem is simply summing the number of edges connecting nodes from different subsets (nodes marked as 0 that are connected to nodes marked as $1$, so the ’cut’ passes through them):
$C(\vec{x})=\sum_{i,j\in E(G)}x_{i}(1-x_{j})+(1-x_{i})x_{j}$ (17)
The Max-Cut problem is equivalent to minimizing an Ising model Hamiltonian, as we now show.
In Ising model Hamiltonians defined over a graph $G=(V,E)$, the Hamiltonian is of the form:
$\mathbb{H}=-\sum_{i,j\in E(G)}J_{ij}s_{i}s_{j}$ (18)
Here each vertex $i$ of the graph is a spin site that can take a spin value $s_{i}=\pm 1$. A spin configuration partitions $V(G)$ into two sets, those with spin up $V^{+}$ and those with spin down $V^{-}$. We use $\delta(V^{+})$ to denote the set of edges that connect the two sets. We can then rewrite the Hamiltonian as:
$\begin{split}\mathbb{H}=-\sum_{i,j\in E(V^{+})}J_{ij}-\sum_{i,j\in E(V^{-})}J_{ij}+\sum_{i,j\in\delta(V^{+})}J_{ij}\\ =-\sum_{i,j\in E(G)}J_{ij}+2\sum_{i,j\in\delta(V^{+})}J_{ij}\\ =C+2\sum_{i,j\in\delta(V^{+})}J_{ij}\end{split}$ (19)
Where $C$ does not depend on the choice of the cut. By setting $J_{ij}=-1$, we get that minimizing this Hamiltonian is equivalent to solving Max-Cut for $G$. Moving from the classical Ising to its quantum counterpart, we get the following equation:
$\mathbb{H}=\sum_{i,j\in E(G)}Z_{i}Z_{j}$ (20)
which can be measured efficiently on a quantum computer by measuring the expectation values of the Pauli $ZZ$ operators. While Max-Cut is a maximization problem, the best assignment for $G$ would translate to the ground state of $\mathbb{H}$ as defined in Eq. 20. Thus, to find the maximum cut for $G$, we seek to minimize the energy of $\mathbb{H}$.